Without actually performing the long division, state whether the following rational numbers will have a terminating decimal expansion or a non-terminating repeating decimal expansion:
35/50 

Grade:12

2 Answers

Harshit Singh
2 years ago
Dear Student

35/50= 7/10
7/10 = 0.7

Thanks
Ram Kushwah
2 years ago
The given number is 35/50
 
Here the Denominator
 
50 = 2 x25
=2^1 x 5^2
 
This is a form of 2^m x 5 ^n where m=1 and n=2
 
Thus as per rules the decimal expansion of 35/50 wii be terminating